# Regularized Q-Learning with Linear Function Approximation

### v1

@@ 10 Proofs of Other Lemmas
$\sum_{t=0}^{T_{\alpha}-1}\mathbb{E}\left\|\omega^{*}(\theta^{t})-\omega^{t+1}\right\|^{2}\leq$ $T_{\alpha}\cdot\frac{4\left(R_{\max}+\delta\right)^{2}}{\lambda_{g}^{2}}$
$=$ $\mathcal{O}\left(\log T\right).$ (35)
Combining (10.3) and (10.3) yields
$\frac{1}{T}\sum_{t=0}^{T-1}$ $\mathbb{E}\left\|\omega^{*}(\theta^{t})-\omega^{t+1}\right\|^{2}$
$\leq$ $\mathcal{O}\left(\frac{\log T}{T}\right)+\frac{8\left(R_{\max}+\delta\right)^{2}}{T^{1/2}\beta_{0}\lambda_{g}^{3}}+\frac{16\alpha_{0}^{2}}{T^{1/2}\beta_{0}^{2}\lambda_{g}^{4}}$
$+\frac{16\alpha_{0}\left(R_{\max}+\delta\right)^{2}}{T^{3/4}\lambda_{g}^{3}}+\frac{8\beta_{0}\left(R_{\max}+\delta\right)^{2}}{T^{1/2}\lambda_{g}^{3}}$
$=$ $\mathcal{O}\left(T^{-1/2}\right).$
10.4 Proof of Lemma 9.2
@@
$J\left(\theta^{t+1}\right)-J\left(\theta^{t}\right)\leq$ $\left\langle\nabla J\left(\theta^{t}\right),\theta^{t+1}-\theta^{t}\right\rangle$
$+\frac{L_{0}+L_{1}\left\|\omega^{*}\left(\theta^{t}\right)-\theta^{t}\right\|}{2}\left\|\theta^{t+1}-\theta^{t}\right\|^{2}$
$\leq$ $\left\langle\nabla J\left(\theta^{t}\right),-\frac{\alpha}{\|\omega^{t+1}-\theta^{t}\|}{h}_{f}^{t}\right\rangle$
$+2\alpha^{2}L_{0}+2\alpha^{2}L_{1}\left\|\omega^{*}\left(\theta^{t}\right)-\theta^{t}\right\|$
$\leq$ $\left\langle\nabla J\left(\theta^{t}\right),-\frac{\alpha}{\|\omega^{t+1}-\theta^{t}\|}\tilde{h}_{f}^{t}\right\rangle$
$+\left\langle\nabla J\left(\theta^{t}\right),-\frac{\alpha}{\|\omega^{t+1}-\theta^{t}\|}\left(h_{f}^{t}-\tilde{h}_{f}^{t}\right)\right\rangle$
@@
where the second inequality follows from Lemma 8.7. We first focus on the first term on the right-hand side (10.4). By taking the conditional expectation on $\mathcal{F}^{\prime}_{t}$, it becomes
$\mathbb{E}\left[\left\langle\nabla J\left(\theta^{t}\right),-\frac{\alpha}{\|\omega^{t+1}-\theta^{t}\|}\tilde{h}_{f}^{t}\right\rangle\mid\mathcal{F}^{\prime}_{t}\right]$
$=$ $\left\langle\nabla J\left(\theta^{t}\right),-\frac{\alpha}{\|\omega^{t+1}-\theta^{t}\|}\bar{\nabla}f(\theta^{t},\omega^{t+1})\right\rangle$
$=$ $\bigg{\langle}\nabla J\left(\theta^{t}\right)-\bar{\nabla}f\left(\theta^{t},\omega^{t+1}\right),-\frac{\alpha}{\|\omega^{t+1}-\theta^{t}\|}\bar{\nabla}f\left(\theta^{t},\omega^{t+1}\right)\bigg{\rangle}$
+¯$+\left\langle\bar{\nabla}f\left(\theta^{t},\omega^{t+1}\right),-\frac{\alpha}{\|\omega^{t+1}-\theta^{t}\|}\bar{\nabla}f\left(\theta^{t},\omega^{t+1}\right)\right\rangle$
$\leq$ $\alpha\left\|\nabla J\left(\theta^{t}\right)-\bar{\nabla}f\left(\theta^{t},\omega^{t+1}\right)\right\|\frac{\left\|\bar{\nabla}f\left(\theta^{t},\omega^{t+1}\right)\right\|}{\|\omega^{t+1}-\theta^{t}\|}$
$-\frac{\alpha}{\|\omega^{t+1}-\theta^{t}\|}\left\|\bar{\nabla}f\left(\theta^{t},\omega^{t+1}\right)\right\|^{2}$
$\leq$ $4\alpha\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{t+1}\right\|-\frac{\alpha\left\|\bar{\nabla}f\left(\theta^{t},\omega^{t+1}\right)\right\|^{2}}{\|\omega^{t+1}-\theta^{t}\|},$ (37)
@@
where the last inequality is due to Lemma 8.9. Under Assumption 3.4, the $\ell_{2}$-norm of the surrogate gradient is lower bounded as
¯$\left\|\bar{\nabla}f\left(\theta^{t},\omega^{t+1}\right)\right\|\geq\sigma_{\min}\left\|\omega^{t+1}-\theta^{t}\right\|.$ (38)
Thus we can bound (10.4) as
$\mathbb{E}$ $\left[\left\langle\nabla J\left(\theta^{t}\right),-\frac{\alpha}{\|\omega^{t+1}-\theta^{t}\|}\cdot\tilde{h}_{f}^{t}\right\rangle\mid\mathcal{F}^{\prime}_{t}\right]$
$\leq 4\alpha\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{t+1}\right\|-\alpha\sigma_{\min}\left\|\bar{\nabla}f\left(\theta^{t},\omega^{t+1}\right)\right\|.$ (39)
Next, we analyze the second term on the right-hand side in (10.4). By taking the conditional expectation on $\mathcal{F}_{t}^{\prime}$, we obtain
$\mathbb{E}\left[\left\langle\nabla J\left(\theta^{t}\right),-\frac{\alpha}{\|\omega^{t+1}-\theta^{t}\|}\left(h_{f}^{t}-\tilde{h}_{f}^{t}\right)\right\rangle\mid\mathcal{F}^{\prime}_{t}\right]$
$\leq$ $\alpha\left\|\nabla J\left(\theta^{t}\right)\right\|\frac{\left\|\mathbb{E}\left[h_{f}^{t}-\tilde{h}_{f}^{t}\mid\mathcal{F}^{\prime}_{t}\right]\right\|}{\|\omega^{t+1}-\theta^{t}\|}$
$\leq$ $\alpha\left\|\nabla J\left(\theta^{t}\right)\right\|d_{TV}\left(\mathbb{P}\left(\left(s_{t},a_{t},s_{t}^{\prime}\right)\in\cdot\mid s_{0}\sim\mu_{0},\pi_{\mathrm{bhv}},P\right),\mathcal{D}\right)$
×$\times\sup_{s,a,s^{\prime}}\|(\gamma z\left(s^{\prime};\tau,\delta,\theta^{t}\right)\sum_{a^{\prime}\in\mathcal{A}}\pi_{\theta^{t}}\left(a^{\prime}\mid s^{\prime}\right)\phi(s^{\prime},a^{\prime})-\phi(s,a))\|$
$\leq$ $2\alpha\left\|\nabla J\left(\theta^{t}\right)\right\|d_{TV}\left(\mathbb{P}\left(\left(s_{t},a_{t},s_{t}^{\prime}\right)\in\cdot\mid s_{0}\sim\mu_{0},\pi_{\mathrm{bhv}},P\right),\mathcal{D}\right)$
$\leq$ $2\alpha^{2}\left\|\nabla J\left(\theta^{t}\right)\right\|$
$\leq$ $2\alpha^{2}\left\|\bar{\nabla}_{\theta}f\left(\theta^{t},\omega^{t+1}\right)\right\|+4\alpha^{2}\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{t+1}\right\|,$ (40)
where the second inequality follows the definition of the total-variation norm and the fourth inequality uses Lemma 1 in [48]. The last inequality is due to Lemma 8.9.
For the third term in (10.4), we bound it by leveraging Assumption 3.4. Given the triangle inequality, we can derive
$\left\|\omega^{*}\left(\theta^{t}\right)-\theta^{t}\right\|\leq$ $\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{t+1}\right\|+\left\|\theta^{t}-\omega^{t+1}\right\|$
$\leq$ $\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{t+1}\right\|+\frac{\left\|\bar{\nabla}f\left(\theta^{t},\omega^{t+1}\right)\right\|}{\sigma_{\min}},$
where the last inequality follows from (38). Taking full expectation in (10.4) and plugging (10.4), (10.4) and the preceding inequality yield
$\mathbb{E}$ $J\left(\theta^{t+1}\right)-\mathbb{E}J\left(\theta^{t}\right)$
@@
$\leq\alpha\left(\alpha\left(\frac{2L_{1}}{\sigma_{\min}}+2\right)-\sigma_{\min}\right)\mathbb{E}\left\|\bar{\nabla}f\left(\theta^{t},\omega^{t+1}\right)\right\|$
@@
$\ \ \ +2\alpha\left(2+\alpha^{2}L_{1}+2\alpha\right)\mathbb{E}\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{t+1}\right\|+2\alpha^{2}L_{0}.$
@@
Let $\eta\triangleq\sigma_{\min}-\alpha({2L_{1}}/{\sigma_{\min}}+2)$. Given the conditions $\alpha<\frac{\sigma_{\min}^{2}}{2(L_{1}+\sigma_{\min})}$ and $\sigma_{\min}\leq 1$, we can derive
$\mathbb{E}J\left(\theta^{t+1}\right)-\mathbb{E}J\left(\theta^{t}\right)\leq$ $-\alpha\eta\mathbb{E}\left\|\bar{\nabla}f\left(\theta^{t},\omega^{t+1}\right)\right\|$
$+6\alpha\mathbb{E}\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{t+1}\right\|+2\alpha^{2}L_{0}.$ (41)
Summing it from $t=T_{\alpha}$ to $t=T-1$ gives
$\!\!\!\!\!\alpha\eta\sum_{t=T_{\alpha}}^{T-1}\mathbb{E}\left\|\bar{\nabla}f\left(\theta^{t},\omega^{t+1}\right)\right\|$
$\leq$ $\mathbb{E}J\left(\theta^{T_{\alpha}}\right)+6\alpha\sum_{t=T_{\alpha}}^{T-1}\mathbb{E}\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{t+1}\right\|+2\alpha^{2}TL_{0}.$ (42)
To continue, we consider the first term in (10.4). By the triangle inequality and Lemma 8.7, we have
$\left\|\theta^{T_{\alpha}}\right\|\leq$ $\sum_{t=0}^{T_{\alpha}-1}\left\|\theta^{t+1}-\theta^{t}\right\|+\left\|\theta^{0}\right\|$
$\leq$ $2\alpha T_{\alpha}+\left\|\theta^{0}\right\|.$
@@
$\mathbb{E}J\left(\theta^{T_{\alpha}}\right)\leq$ $\left\|\omega^{*}\left(\theta^{T_{\alpha}}\right)-\theta^{T_{\alpha}}\right\|^{2}$
$\leq$ $2\left\|\omega^{*}\left(\theta^{T_{\alpha}}\right)\right\|^{2}+2\left\|\theta^{T_{\alpha}}\right\|^{2}$
$\leq$ $\frac{2\left(R_{\max}+\delta\right)^{2}}{\lambda_{g}^{2}}+16\alpha^{2}T_{\alpha}^{2}+4\left\|\theta^{0}\right\|^{2},$ (43)
@@
For the second term in (10.4), we bound it by using (9.1):
$\sum_{t=T_{\alpha}}^{T-1}\mathbb{E}\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{t+1}\right\|=\mathcal{O}(T^{3/4})$ (44)
Substituting (10.4) and (44) into (10.4) and dividing $\alpha\eta$ on both sides yield that
$\sum_{t=T_{\alpha}}^{T-1}\mathbb{E}\left\|\bar{\nabla}f\left(\theta^{t},\omega^{t+1}\right)\right\|\leq$ $\frac{2\left(R_{\max}+\delta\right)^{2}}{\alpha\eta\lambda_{g}^{2}}+\frac{16\alpha T_{\alpha}^{2}}{\eta}+\frac{4\left\|\theta^{0}\right\|^{2}}{\alpha\eta}$
$+\frac{2\alpha TL_{0}}{\eta}+\mathcal{O}(T^{3/4}).$ (45)
For $t<T_{\alpha}$, we can bound the $\ell_{2}$-norm of the surrogate gradient as
@@
$\leq$ $2\left(\frac{R_{\max}+\delta}{\lambda_{g}}+2\alpha T_{\alpha}+\left\|\theta^{0}\right\|\right),$ (46)
@@
where the last inequality follows from Lemma 8.7. As a result, we have
@@
$\sum_{t=0}^{T_{\alpha}-1}\left\|\bar{\nabla}f\left(\theta^{t},\omega^{t+1}\right)\right\|\leq$ $2\left(\frac{R_{\max}+\delta}{\lambda_{g}}+2\alpha T_{\alpha}+\left\|\theta^{0}\right\|\right)T_{\alpha}$
$=$ $\mathcal{O}\left(\log T\right).$ (47)
Combining (10.4) and (10.4), and dividing $T$ lead to

### v3

@@ C.4 Proof of Lemma B1
$\sum_{t=0}^{T_{\alpha}-1}\mathbb{E}\left\|\omega^{*}(\theta^{t})-\omega^{t+1}\right\|^{2}\leq T_{\alpha}\cdot\frac{4\left(R_{\max}+\delta\right)^{2}}{\lambda_{g}^{2}}=\mathcal{O}\left(\log T\right).$ (36)
Combining (C.4) and (36) yields
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left\|\omega^{*}(\theta^{t})-\omega^{t+1}\right\|^{2}\leq$ $\mathcal{O}\left(\frac{\log T}{T}+\frac{1}{\beta T}+\frac{\alpha^{2}}{\beta^{2}}+\beta\log T+\alpha+\beta+\alpha\log T\right)$
$=$ $\mathcal{O}\left(\frac{\log T}{T^{1/2}}\right),$
where the last equality is given by $\alpha=\alpha T^{-3/4}$ and $\beta=\beta T^{-1/2}$.
C.5 Proof of Lemma B2
Define
$o^{t}_{f}(\theta,\omega)\triangleq\bigg{(}\gamma z\left(s^{\prime}_{t};\tau,\delta,\theta\right)\sum_{a^{\prime}\in\mathcal{A}}\pi_{\theta}\left(a^{\prime}\mid s^{\prime}_{t}\right)\phi(s_{t}^{\prime},a^{\prime})-\phi(s_{t},a_{t})\bigg{)}\phi(s_{t},a_{t})^{\top}\left(\omega-\theta\right),$
and
$\xi_{f}^{t}\left(\theta,\omega\right)\triangleq\left\langle\bar{\nabla}_{\theta}f(\theta,\omega),\frac{1}{\|\omega-\theta\|}\left(\bar{\nabla}_{\theta}f(\theta,\omega)-o^{t}_{f}(\theta,\omega)\right)\right\rangle.$
$\xi^{t}_{g}$ has the following properties.
Lemma C2.
For $t\geq 0$,
  1. 1.
    $\xi_{f}^{t}(\theta,\omega)$ is $24$-Lipschitz in $\omega$ for $\omega\in\Omega$;
  2. 2.
    for $\theta_{1},\theta_{2}\in\mathbb{R}^{d},\omega\in\Omega$,
    $\left|\xi^{t}_{f}(\theta_{1},\omega)-\xi^{t}_{f}(\theta_{2},\omega)\right|\leq\left(\frac{8L_{1}}{\sigma_{\min}}\|\bar{\nabla}_{\theta}f(\theta_{1},\omega)\|+24\right)\left\|\theta_{1}-\theta_{2}\right\|.$
The proof for Lemma C2 is in Appendix C.7.
@@ C.5 Proof of Lemma B2
$J\left(\theta^{t+1}\right)-J\left(\theta^{t}\right)\leq$ $\left\langle\nabla J\left(\theta^{t}\right),\theta^{t+1}-\theta^{t}\right\rangle+\frac{L_{0}+L_{1}\left\|\omega^{*}\left(\theta^{t}\right)-\theta^{t}\right\|}{2}\left\|\theta^{t+1}-\theta^{t}\right\|^{2}$
$\leq$ $\left\langle\nabla J\left(\theta^{t}\right),-\frac{\alpha}{\|\omega^{t+1}-\theta^{t}\|}{h}_{f}^{t}\right\rangle+2\alpha^{2}L_{0}+2\alpha^{2}L_{1}\left\|\omega^{*}\left(\theta^{t}\right)-\theta^{t}\right\|$
$\leq$ $-\alpha\frac{\left\|\bar{\nabla}_{\theta}f(\theta^{t},\omega^{t+1})\right\|^{2}}{\left\|\omega^{t+1}-\theta^{t}\right\|}+\left\langle\nabla J\left(\theta^{t}\right)-\bar{\nabla}_{\theta}f(\theta^{t},\omega^{t+1}),-\frac{\alpha}{\|\omega^{t+1}-\theta^{t}\|}h_{f}^{t}\right\rangle$
@@
where the second inequality follows from Lemma A4. Under Assumption 3.2, the $\ell_{2}$-norm of the surrogate gradient is lower bounded as
¯$\left\|\bar{\nabla}_{\theta}f\left(\theta^{t},\omega^{t+1}\right)\right\|\geq\left\|\omega^{t+1}-\theta^{t}\right\|.$ (38)
Thus we can bound the first term on the right-hand side of (C.5) as
@@
$-\alpha\frac{\left\|\bar{\nabla}_{\theta}f(\theta^{t},\omega^{t+1})\right\|^{2}}{\left\|\omega^{t+1}-\theta^{t}\right\|}\leq-\alpha\sigma_{\min}\left\|\bar{\nabla}_{\theta}f(\theta^{t},\omega^{t+1})\right\|.$ (39)
For the second term on the right-hand side of (C.5),
@@
$\left\langle\nabla J\left(\theta^{t}\right)-\bar{\nabla}_{\theta}f(\theta^{t},\omega^{t+1}),-\frac{\alpha}{\|\omega^{t+1}-\theta^{t}\|}h_{f}^{t}\right\rangle$
$\leq$ $\alpha\left\|\nabla J\left(\theta^{t}\right)-\bar{\nabla}_{\theta}f\left(\theta^{t},\omega^{t+1}\right)\right\|\frac{\left\|h_{f}^{t}\right\|}{\|\omega^{t+1}-\theta^{t}\|}$
$\leq$ $4\alpha\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{t+1}\right\|,$ (40)
where the last inequality is due to Lemma A5.
Next, we focus on the third term on the right-hand side of (C.5). We define
$\tilde{\xi}_{f}(\theta,\omega)=\bigg{\langle}\frac{\bar{\nabla}_{\theta}f(\theta,\omega)}{\|\omega-\theta\|},\Big{(}\bar{\nabla}_{\theta}f(\theta,\omega)-\gamma z\left(\tilde{s}^{\prime};\tau,\delta,\theta\right)\big{(}\sum_{a^{\prime}\in\mathcal{A}}\pi_{\theta}\left(a^{\prime}\mid\tilde{s}^{\prime}\right)\phi(\tilde{s}^{\prime},a^{\prime})-\phi(\tilde{s},\tilde{a})\big{)}\phi(\tilde{s},\tilde{a})^{\top}\left(\omega-\theta\right)\Big{)}\bigg{\rangle},$
where $(\tilde{s},\tilde{a},\tilde{s}^{\prime})\sim\mathcal{D}$. For $t\geq T_{\beta}+1$, by taking the conditional expectation on $\mathcal{F}_{t}^{\prime}$, we obtain
$\mathbb{E}\left[\xi^{t}_{f}(\theta^{t},\omega^{t+1})\mid\mathcal{F}^{\prime}_{t-T_{\beta}-1}\right]$
$\leq$ $\mathbb{E}\left[\xi^{t}_{f}(\theta^{t},\omega^{t+1})-\xi^{t}_{f}(\theta^{t-T_{\beta}-1},\omega^{t+1})\mid\mathcal{F}^{\prime}_{t-T_{\beta}-1}\right]$
$+\mathbb{E}\left[\xi^{t}_{f}(\theta^{t-T_{\beta}-1},\omega^{t+1})-\xi^{t}_{f}(\theta^{t-T_{\beta}-1},\omega^{t-T_{\beta}})\mid\mathcal{F}^{\prime}_{t-T_{\beta}-1}\right]$
$+\mathbb{E}\left[\xi^{t}_{f}(\theta^{t-T_{\beta}-1},\omega^{t-T_{\beta}})-\tilde{\xi}_{f}(\theta^{t-T_{\beta}-1},\omega^{t-T_{\beta}})\mid\mathcal{F}^{\prime}_{t-T_{\beta}-1}\right]$
$\leq$ $\alpha(T_{\beta}+1)\left(\frac{16L_{1}}{\sigma_{\min}}\mathbb{E}\left[\|\bar{\nabla}_{\theta}f(\theta^{t},\omega^{t+1})\|\mid\mathcal{F}^{\prime}_{t-T_{\beta}-1}\right]+48\right)$
$+48\beta(T_{\beta}+1)(R_{\max}+\delta)/\lambda_{g}$
$+2\left\|\bar{\nabla}_{\theta}f(\theta^{t-T_{\beta}-1},\omega^{t-T_{\beta}})\right\|d_{TV}\left(\mathbb{P}\left(\left(s_{t},a_{t},s_{t}^{\prime}\right)\in\cdot\mid s_{t-T_{\beta}},\pi_{\mathrm{bhv}},P\right),\mathcal{D}\right)$
$\leq$ $\alpha(T_{\beta}+1)\left(\frac{16L_{1}}{\sigma_{\min}}\mathbb{E}\left[\|\bar{\nabla}_{\theta}f(\theta^{t},\omega^{t+1})\|\mid\mathcal{F}^{\prime}_{t-T_{\beta}-1}\right]+48\right)$
$+48\beta(T_{\beta}+1)(R_{\max}+\delta)/\lambda_{g}$
$+2\beta\mathbb{E}\left[\left\|\bar{\nabla}_{\theta}f(\theta^{t-T_{\beta}-1},\omega^{t-T_{\beta}})-\bar{\nabla}_{\theta}f(\theta^{t-T_{\beta}-1},\omega^{t+1})\right\|\right]$
$+2\beta\mathbb{E}\left[\left\|\bar{\nabla}_{\theta}f(\theta^{t-T_{\beta}-1},\omega^{t+1})-\bar{\nabla}_{\theta}f(\theta^{t},\omega^{t+1})\right\|\mid\mathcal{F}^{\prime}_{t-T_{\beta}-1}\right]$
$+2\beta\mathbb{E}\left[\left\|\bar{\nabla}_{\theta}f(\theta^{t},\omega^{t+1})\right\|\mid\mathcal{F}^{\prime}_{t-T_{\beta}-1}\right]$
$\leq$ $\left((16+4\beta)\alpha(T_{\beta}+1)\frac{L_{1}}{\sigma_{\min}}+2\beta\right)\mathbb{E}\left[\left\|\bar{\nabla}_{\theta}f(\theta^{t},\omega^{t+1})\right\|\mid\mathcal{F}^{\prime}_{t-T_{\beta}-1}\right]$
$+(48+8\beta)\alpha(T_{\beta}+1)+48\beta(T_{\beta}+1)(R_{\max}+\delta)/\lambda_{g}.$ (41)
The last term in (C.5) can be bounded as
$2\alpha^{2}L_{1}\left\|\omega^{*}(\theta^{t})-\theta^{t}\right\|$
$\leq$ $2\alpha^{2}L_{1}\left\|\omega^{*}(\theta^{t})-\omega^{t+1}\right\|+2\alpha^{2}L_{1}\left\|\theta^{t}-\omega^{t+1}\right\|$
$\leq$ $2\alpha^{2}L_{1}\left\|\omega^{*}(\theta^{t})-\omega^{t+1}\right\|+2\alpha^{2}L_{1}\frac{\left\|\bar{\nabla}_{\theta}f(\theta^{t},\omega^{t+1})\right\|}{\sigma_{\min}},$
where the last inequality follows from (38).
Taking full expectation in (C.5) and plugging (39), (C.5), (C.5) and the preceding inequality yield
$\mathbb{E}\left[J(\theta^{t+1})-J(\theta^{t})\right]$
$\leq$ $-\alpha(\sigma_{\min}-((16\alpha(T_{\beta}+1)+4\alpha\beta(T_{\beta}+1)+2\alpha^{2})\frac{L_{1}}{\sigma_{\min}}+2\beta))$
×$\times\mathbb{E}\left[\left\|\bar{\nabla}_{\theta}f(\theta^{t},\omega^{t+1})\right\|\right]$
$+(4\alpha+2\alpha^{2}L_{1})\mathbb{E}\left\|\omega^{*}(\theta^{t})-\omega^{t+1}\right\|+48\alpha^{2}(T_{\beta}+1)$
$+8\alpha^{2}\beta(T_{\beta}+1)+48\alpha\beta(T_{\beta}+1)(R_{\max}+\delta)/\lambda_{g}+2\alpha^{2}L_{0}.$
@@
Given the conditions
$\alpha\leq\frac{\beta}{\frac{\log(\beta/\kappa)}{\log\rho}+2},\beta\leq$ $\frac{\sigma_{\min}}{4},$
$64\beta+16\beta^{2}+8\alpha^{2}\leq$ $\frac{\sigma_{\min}^{2}}{L_{1}},$
we have
$\mathbb{E}\left[J(\theta^{t+1})-J(\theta^{t})\right]$
$\leq$ $-\frac{\alpha\sigma_{\min}}{4}\mathbb{E}\left[\left\|\bar{\nabla}_{\theta}f(\theta^{t},\omega^{t+1})\right\|\right]+(4\alpha+2\alpha^{2}L_{1})\mathbb{E}\left\|\omega^{*}(\theta^{t})-\omega^{t+1}\right\|$
$+48\alpha^{2}(T_{\beta}+1)+8\alpha^{2}\beta(T_{\beta}+1)$
$+48\alpha\beta(T_{\beta}+1)(R_{\max}+\delta)/\lambda_{g}+2\alpha^{2}L_{0}.$
Summing this inequality from $t=T_{\beta}+1$ to $t=T-1$ gives:
$\frac{\alpha\sigma_{\min}}{4}\sum_{t=T_{\beta}+1}^{T-1}\mathbb{E}\left[\left\|\bar{\nabla}_{\theta}f(\theta^{t},\omega^{t+1})\right\|\right]$
$\leq$ $\mathbb{E}\left[J(\theta^{T_{\beta}+1})\right]+(4\alpha+2\alpha^{2}L_{1})\sum_{t=T_{\beta}+1}^{T-1}\mathbb{E}\left\|\omega^{*}(\theta^{t})-\omega^{t+1}\right\|$
$+48\alpha^{2}(T_{\beta}+1)T+8\alpha^{2}\beta(T_{\beta}+1)T$
$+48\alpha\beta(T_{\beta}+1)T(R_{\max}+\delta)/\lambda_{g}+2\alpha^{2}TL_{0}.$ (42)
To continue, we consider the first term in (C.5). By the triangle inequality and Lemma A4, we have
$\left\|\theta^{T_{\beta}+1}\right\|\leq\sum_{t=0}^{T_{\beta}}\left\|\theta^{t+1}-\theta^{t}\right\|+\left\|\theta^{0}\right\|\leq 2\alpha(T_{\beta}+1)+\left\|\theta^{0}\right\|.$
@@
$\mathbb{E}J\left(\theta^{T_{\beta}+1}\right)$
$\leq$ $\left\|\omega^{*}\left(\theta^{T_{\beta}+1}\right)-\theta^{T_{\beta}+1}\right\|^{2}$
$\leq$ $2\left\|\omega^{*}\left(\theta^{T_{\beta}+1}\right)\right\|^{2}+2\left\|\theta^{T_{\beta}+1}\right\|^{2}$
$\leq$ $\frac{2\left(R_{\max}+\delta\right)^{2}}{\lambda_{g}^{2}}+16\alpha^{2}(T_{\beta}+1)^{2}+4\left\|\theta^{0}\right\|^{2},$ (43)
@@
For the second term in (C.5), we bound it by using (21):
$\sum_{t=T_{\beta}+1}^{T-1}\mathbb{E}\left\|\omega^{*}\left(\theta^{t}\right)-\omega^{t+1}\right\|=\mathcal{O}(T^{3/4}(\log T)^{1/2}).$ (44)
Substituting (C.5) and (44) into (C.5) and dividing $\alpha\sigma_{\min}/4$ on both sides yield that
$\sum_{t=T_{\beta}+1}^{T-1}\mathbb{E}\left\|\bar{\nabla}_{\theta}f\left(\theta^{t},\omega^{t+1}\right)\right\|$
$\leq$ $\frac{8\left(R_{\max}+\delta\right)^{2}}{\alpha\sigma_{\min}\lambda_{g}^{2}}+\frac{64\alpha(T_{\beta}+1)^{2}}{\sigma_{\min}}+\frac{16\left\|\theta^{0}\right\|^{2}}{\alpha\sigma_{\min}}$
$+\frac{192\alpha(T_{\beta}+1)T}{\sigma_{\min}}+\frac{32\alpha\beta(T_{\beta}+1)T}{\sigma_{\min}}$
$+\frac{192\beta(T_{\beta}+1)T(R_{\max}+\delta)}{\lambda_{g}\sigma_{\min}}$
$+\frac{8\alpha TL_{0}}{\sigma_{\min}}+\mathcal{O}(T^{3/4}(\log T)^{1/2}).$ (45)
For $t\leq T_{\beta}$, we can bound the $\ell_{2}$-norm of the surrogate gradient as
@@
$\leq$ $2\left(\frac{R_{\max}+\delta}{\lambda_{g}}+2\alpha T_{\beta}+\left\|\theta^{0}\right\|\right),$
@@
where the last inequality follows from Lemma A4. As a result, we have
@@
$\sum_{t=0}^{T_{\beta}-1}\left\|\bar{\nabla}f\left(\theta^{t},\omega^{t+1}\right)\right\|\leq 2\left(\frac{R_{\max}+\delta}{\lambda_{g}}+2\alpha T_{\beta}+\left\|\theta^{0}\right\|\right)T_{\beta}=\mathcal{O}\left(\log T\right).$ (46)
Combining (C.5) and (46), and dividing $T$ lead to